Find the unit rate in feet per second.
180 Miles
-------------
2 Hours

Found 2 solutions by Cromlix, stanbon:
Answer by Cromlix(4381)

Hi there,
180 miles in 2 hours
180 miles = 950400 ft.
2 hours = 7,200 secs.
950400 ft in 7,200 secs
950400/7,200 = 132 ft per second
Hope this helps :-)

Answer by stanbon(75887)

Find the unit rate in feet per second.
180 miles/2 hours = (90 miles/hour)*(5280 ft/1 mile)*(1 hour/3600 sec)
-------
= (90*5280*1)/(1*1*3600)(ft/sec)
------
= 132 (ft/sec)
---------
Cheers,
Stan H.
-----------